OpenStudy (df001):

Find solution in (0,2pi) (2/3)Cot(x)-(5/6)=-(3/2)

OpenStudy (df001):

\[\frac{ 2 }{ 3 }Cotx-\frac{ 5 }{ 6 }=-\frac{ 3 }{ 2 }\]

OpenStudy (df001):

plz help me i need this.. i have 8 questions on homework and i cant even solve one of them.. im in stress

OpenStudy (campbell_st):

ok.. so add 5/6 to both sides of the equation... \[\frac{2\cot(x)}{3} = \frac{14}{6}\] next mutliply by 3 and then divide by to... or multiply both sides of the equation by 3/2 \[\cot(x) = \frac{14}{6} \times \frac{3}{2}\] that should help

OpenStudy (df001):

the topic for this is trigonometric equation

OpenStudy (campbell_st):

I can see that... does the value of x have a stated domain...?

OpenStudy (df001):